# CMRNext: Camera to LiDAR Matching in the Wild for Localization and Extrinsic Calibration

### v4

@@ I Introduction
Localization is an essential task for any autonomous robot, as it is a precursor for subsequent safety-critical tasks such as path planning and navigation. The accuracy of Global Navigation Satellite Systems is not sufficient for most of these tasks, and therefore, sensors such as Light Detection And Ranging (LiDAR) and cameras are employed to estimate the pose of the robot. While LiDAR-based localization methods [1, 2] achieve the best performance in terms of accuracy and robustness, the comparatively high price of LiDAR sensors hinders their widespread adoption. On the other hand, cameras are inexpensive and already widely available in consumer vehicles for Advanced Driver Assistance Systems (ADAS) applications. However, the performance of camera-based localization methods [3, 4, 5, 6] is still not comparable to LiDAR-based localization techniques, especially when the lighting and weather conditions during deployment differ from those recorded while mapping.
@@
Consequently, monocular localization in LiDAR maps has gained considerable attention from both academia and industry in recent years, as it combines the high accuracy and robustness of LiDAR-based localization while only requiring a camera sensor during deployment. The increasing availability of LiDAR maps from mapping companies such as HERE and TomTom has further contributed to the popularity of this approach. Once a LiDAR map is available, either obtained from a mapping company or generated using recent LiDAR Simultaneouos Localization And Mapping (SLAM) algorithms [7, 2, 8], a vehicle can be localized within this map using a single camera. Considering the low cost of cameras, compared to LiDAR sensors, this approach is particularly appealing for large-scale deployment, where the cost of the final system is a key factor. While traditionally this task has been addressed by first generating a 3D representation from the camera image and then matching it against the LiDAR map [9], the recent advancement of Deep Neural Network (DNN)-based methods for multimodal representation learning [10, 11, 12] has enabled the matching to be performed directly in the feature space.
@@
Most of the existing DNN-based approaches for monocular localization in LiDAR maps, however, follow the end-to-end learning paradigm, where the network directly predicts the metric pose of the camera given an input image and a representation of the LiDAR map. While this paradigm has the advantage of being simple to train and usually achieves good performance when evaluated on images from the same camera used during training, it has the major drawback of being dependent on the specific camera’s intrinsic parameters. Consequently, the network has to be retrained for every new camera and cannot generalize to different sensor setups. This is even worse when these approaches are applied to the task of extrinsic calibration, where the goal is to estimate the relative pose between the camera and the LiDAR sensor. In this case, end-to-end approaches trained on a single camera-LiDAR pair typically tend to overfit to the specific intrinsic and extrinsic parameters of the dataset and fail to generalize even when evaluated on a different camera setup in the same dataset. For example, a network trained in an end-to-end manner on the left camera of the KITTI dataset [13] and evaluated on the right camera will still predict the pose of the left camera. This shows that these networks just learn the specific extrinsic parameters of the camera and thus cannot be used for extrinsic calibration in any real-world application.
@@
We propose to leverage uncertainty estimation for camera-LiDAR matches, such that only the most reliable matches are used for the pose estimation step.
@@ II Related Works
The extrinsic calibration between LiDAR and camera sensors is typically performed using target-based methods, where an ad-hoc calibration board is placed in the environment, and the relative pose between the two sensors is estimated by detecting the board in both modalities. For example, common target boards used for this task are one or multiple checkerboards on planar boards [29, 30], ordinary boxes [31] or a planar board with circular holes [32]. Although target-based methods are widely used, they require a data recording specifically for this purpose, which should be repeated periodically to account for possible changes in the extrinsic parameters due to vibrations, temperature variations, and manual repositioning of the sensors. Early attempts at targetless extrinsic calibration used visual similarity measures, such as mutual information [33, 17], contours [34], or edges [35], to estimate the relative pose between the sensors. These methods, however, are sensitive to appearance changes, such as lighting and weather conditions, and thus require a careful tuning of the parameters.
@@ III-A Matching Step
The main intuition behind the matching step is that the projection of the LiDAR-map in the correct pose should align with the camera image, see Figure 3 (top). Therefore, given a camera image and a (misaligned) LiDAR projection, we can train a CNN to align the two inputs by predicting pixel-level displacements. Specifically, we generate a synthesized depth image, which we refer to as a LiDAR-image, by projecting the map into a virtual image plane placed at an initial pose estimate . This initial pose can be obtained, for example, from a GNSS or a cross-modal place recognition method [45]. Although CMRNext is independent of the camera intrinsic parameters, i.e. the network’s weights are not tied to a specific camera, the intrinsics are still required as input to our method, and they are used to perform the projection. Figure 2(b) shows an example of a LiDAR-image.
@@
The architecture of CMRNext is based on RAFT [40], a state-of-the-art network for optical flow estimation between two consecutive RGB frames. More details regarding the network architecture are reported in Section III-C. The output of CMRNext is a dense feature map, which consists of four channels that represent, for every pixel in the LiDAR-image, the displacement (, ) of the pixel in the RGB image from the same world point and the uncertainties of this match (, ). A visual representation of this pixel displacement is depicted in Figure 2 (c). The predicted uncertainties can be used for any downstream task, such as active SLAM [47] or map change detection [48].
@@ III-C Network Architecture
where is the concatenation of the patch from the correlation volume, the output of the context encoder, and the current predicted flow . The operation denotes a convolutional layer with a kernel size of , and and are the weights of the convolutional layers. Afterward, the hidden state is used to predict the next residual flow and the respective uncertainty using two convolutional layers. The residual flow is then added to the previous flow to obtain the updated flow . This process is repeated times to obtain the final flow. Since the GRU weights and are shared across iterations, the number of update iterations can be increased during inference for improved performance. Finally, the predicted flow at every iteration is upsampled to the original image resolution using learned convex upsampling [51].
@@ IV-G Monocular Localization
Initial Pose | Ground Truth | CMRNext (ours) |
IV-G1 Iterative Refinement
IV-H Ablation Study
In Table IV we evaluate the performance of CMRNext using different optical flow networks. In particular, we evaluate established optical flow networks, namely PWCNet [40], MaskFlowNet [66] and RAFT [51], as well as the recent state-of-the-art methods GMFlow [52] and FlowFormer [67]. Unsurprisingly, RAFT outperforms previous methods PWCNet and MaskFlowNet by a large margin. However, we noticed that both transformer-based methods GMFlow and FlowFormer underperformed compared to RAFT, although they achieved better results on optical flow benchmarks. We hypothesize that this is due to the fact that networks based on Vision Transformers (ViT) lack intrinsic inductive bias [68] and thus require a large amount of data to train. The vast majority of methods for optical flow estimation are pretrained on large-scale synthetic datasets, such as FlyingChairs and FlyingThings, and therefore ViT-based methods are able to learn this inductive bias. For our cross-modal task, however, such large-scale datasets are not available, and therefore, the CNN-based method RAFT achieves the best performance among all the evaluated networks. Moreover, the positional encoding used in transformers is sensitive to the shape of the input, and during inference, images with the same resolution as the ones used during training should be used for optimal performance. The authors of FlowFormer [67] suggested a tiling strategy where the input image is split into multiple tiles, each tile is then independently processed by the network, and the resulting optical flow fields are combined together. While this strategy substantially improves the performance of FlowFormer (see Table IV), it requires multiple forward passes of the network, drastically increasing the computational cost.
In Table V, we additionally investigate the impact of the positional encoding presented in Equation 10, the occlusion filter discussed in Section III-A, the voxel size used to discretize the map, and which input to use for the context encoder (camera image or LiDAR image). We first evaluate the effect of the positional encoding by comparing the performance of CMRNext using no positional encoding, the positional encoding presented in Equation 10 with six Fourier frequencies, and with 12 frequencies. The first set of rows in Table V shows that the positional encoding substantially improves the performance of CMRNext, with the best results achieved using 12 frequencies, demonstrating that Fourier features are better suited for deep neural network processing than (possibly normalized) coordinates.
@@ IV-H Ablation Study
Subsequently, we evaluated three different resolutions for the LiDAR maps: , , and . As shown in the third set of rows in Table V, the voxel size has a significant impact on the performance of our method. This result is not surprising, as a coarse resolution would lead to sparse LiDAR-images with non-well-defined edges, hindering the pixel-wise matching with the camera images. Although a resolution lower than could further improve the performance, computational and storage requirements would increase significantly. Therefore, we chose as the default resolution for the LiDAR maps, as a good trade-off between performance and computational cost.
Finally, we compared the performance of CMRNext using the camera image and the LiDAR image as the input for the context encoder. The last two rows of Table V show that the LiDAR context encoder achieves better performance than the camera context encoder.
IV-I LiDAR-Camera Calibration
@@
We compared the performance of CMRNext on the LiDAR-camera extrinsic calibration task with our previous methods CMRNet [21], as well as with the state-of-the-art methods RGGNet [38] and LCCNet [39]. For the latter methods, we used the pretrained models provided by the respective authors444https://github.com/LvXudong-HIT/LCCNet555https://github.com/KleinYuan/RGGNet. Other existing methods [64, 37, 44, 62] do not provide a publicly available implementation and therefore we were unable to evaluate them on the right camera of the KITTI dataset, as discussed above. In Table VI, we report the results of CMRNet, LCCNet, and CMRNext after the iterative refinement (five iterations for CMRNet and LCCNet, and three iterations for CMRNext). The results show that although LCCNet achieves the lowest translation error when evaluated on the same camera used for training, it fails to generalize to a different camera, still predicting the extrinsic for the left camera when evaluated on the right one. We observe a similar behavior for our previous work CMRNet, as both CMRNet and LCCNet are camera-dependent and thus more prone to overfitting to the specific extrinsic. Our proposed CMRNext, on the other hand, outperforms existing methods on the right camera, with a median translation error of and a median rotation error of , while at the same time achieving the lowest rotation error and comparable translation error on the left camera.
@@
where is the initial error applied to the -th sample, and is the left-invariant Riemannian metric on the SE(3) Lie group between the ground truth and the predicted extrinsic calibration. The results reported in Table VII show that CMRNext outperforms both -RegNet and RGGNet on all testing splits by a large margin. When evaluated on the right camera, CMRNext achieves an MMR of 90.13% on the T1 testing split, compared to 12.37% achieved by RGGNet and 89.71% MRR on the T3 split, compared to -113.67% achieved by RGGNet.
@@
IV-J Extrinsic Calibration — Temporal Aggregation
@@
The multi-frames results reported in Table VIII show that the performance of all tendency measures decreases as the number of frames used for aggregation decreases, as hypothesized. Moreover, the mode consistently outperforms the mean measures in all experiments. The median aggregation achieves the best translation error on the left camera but decreases significantly on the right camera, where the mode achieves the best results. When using all frames of the testing sequence (4541 frames), the aggregated mode extrinsic calibration achieves a translation error of and a rotation error of on the left camera, and and on the right camera, setting the new state of the art for targetless LiDAR-camera extrinsic calibration.
IV-K Extrinsic Calibration — Generalization
We evaluate the generalization ability of CMRNext on the extrinsic calibration task on our three in-house platforms detailed in Section IV-A: a self-driving perception car, a quadruped robot, and a UAV. To quantitatively evaluate our proposed method, we first calibrated the extrinsic parameters of each platform using a simple manual calibration procedure. In particular, we recorded a calibration sequence by moving a planar calibration board in front of the platform and then we manually selected the corners of the calibration board in both the camera images and the LiDAR scans. Finally, based on these correspondences, we used PnP with RANSAC, followed by a Levenberg-Marquardt optimization based on the inliers of RANSAC. We compare the translation and rotation errors of CMRNext on the three platforms against the manual calibration. In this experiment, since the number of frames in each dataset is relatively low (around 20 frames), we use 30 different random initial noise poses for each frame, and we report the median frame-wise errors and the error of the pose computed using temporal aggregation across all frames. The results reported in Table IX show that our method with temporal aggregation achieves rotation errors lower than on all three platforms and translation errors lower than on the car and UAV platforms. The translation error on the quadruped is considerably higher than those on the other platforms. However, this higher error might derive from an inaccurate manual calibration, as we have no way to guarantee the accuracy of our manual calibration procedure.

### v5

@@ I Introduction
Localization is an essential task for any autonomous robot, as it is a precursor for subsequent safety-critical tasks such as path planning and navigation. The accuracy of Global Navigation Satellite Systems is not sufficient for most of these tasks, and therefore, sensors such as Light Detection And Ranging (LiDAR) and cameras are employed to estimate the pose of the robot. While LiDAR-based localization methods [1, 2] achieve the best performance in terms of accuracy and robustness, the comparatively high price of LiDAR sensors hinders their widespread adoption. On the other hand, cameras are inexpensive and already widely available in consumer vehicles for Advanced Driver Assistance Systems (ADAS) applications. However, the performance of camera-based localization methods [3, 4, 5, 6, 7] is still not comparable to LiDAR-based localization techniques, especially when the lighting and weather conditions during deployment differ from those recorded while mapping.
@@
Consequently, monocular localization in LiDAR maps has gained considerable attention from both academia and industry in recent years, as it combines the high accuracy and robustness of LiDAR-based localization while only requiring a camera sensor during deployment. The increasing availability of LiDAR maps from mapping companies such as HERE and TomTom has further contributed to the popularity of this approach. Once a LiDAR map is available, either obtained from a mapping company or generated using recent LiDAR Simultaneouos Localization And Mapping (SLAM) algorithms [8, 2, 9], a vehicle can be localized within this map using a single camera. Considering the low cost of cameras, compared to LiDAR sensors, this approach is particularly appealing for large-scale deployment, where the cost of the final system is a key factor. While traditionally this task has been addressed by first generating a 3D representation from the camera image and then matching it against the LiDAR map [10], the recent advancement of Deep Neural Network (DNN)-based methods for multimodal representation learning [11, 12, 13, 14] has enabled the matching to be performed directly in the feature space.
@@
Most of the existing DNN-based approaches for monocular localization in LiDAR maps, however, follow the end-to-end learning paradigm, where the network directly predicts the metric pose of the camera given an input image and a representation of the LiDAR map. While this paradigm has the advantage of being simple to train and usually achieves good performance when evaluated on images from the same camera used during training, it has the major drawback of being dependent on the specific camera’s intrinsic parameters. Consequently, the network has to be retrained for every new camera and cannot generalize to different sensor setups. This is even worse when these approaches are applied to the task of extrinsic calibration, where the goal is to estimate the relative pose between the camera and the LiDAR sensor. In this case, end-to-end approaches trained on a single camera-LiDAR pair typically tend to overfit to the specific intrinsic and extrinsic parameters of the dataset and fail to generalize even when evaluated on a different camera in the same dataset. For example, a network trained in an end-to-end manner on the left camera of the KITTI dataset [15] and evaluated on the right camera will still predict the pose of the left camera. This shows that these networks just learn the specific extrinsic parameters of the camera and thus cannot be used for extrinsic calibration in any real-world application, suggesting that a combination of learning based-methods and geometric techniques should be preferred [16].
@@
We propose to estimate the uncertainty of camera-LiDAR matches as additional output for usage in downstream tasks.
@@ II Related Works
The extrinsic calibration between LiDAR and camera sensors is typically performed using target-based methods, where an ad-hoc calibration board is placed in the environment, and the relative pose between the two sensors is estimated by detecting the board in both modalities. For example, common target boards used for this task are one or multiple checkerboards on planar boards [32], ordinary boxes [33] or a planar board with circular holes [34]. Although target-based methods are widely used, they require a data recording specifically for this purpose, which should be repeated periodically to account for possible changes in the extrinsic parameters due to vibrations, temperature variations, and manual repositioning of the sensors. Early attempts at targetless extrinsic calibration used visual similarity measures, such as mutual information [20], contours [35], or edges [36], to estimate the relative pose between the sensors. These methods, however, are sensitive to appearance changes, such as lighting and weather conditions, and thus require a careful tuning of the parameters.
@@ III-A Matching Step
The main intuition behind the matching step is that the projection of the LiDAR-map in the correct pose should align with the camera image, see Figure 3 (middle). Therefore, given a camera image and a (misaligned) LiDAR projection, we can train a CNN to align the two inputs by predicting pixel-level displacements. Specifically, we generate a synthesized depth image, which we refer to as a LiDAR-image, by projecting the map into a virtual image plane placed at an initial pose estimate . This initial pose can be obtained, for example, from a GNSS or a cross-modal place recognition method [46]. Although CMRNext is independent of the camera intrinsic parameters, i.e. the network’s weights are not tied to a specific camera, the intrinsics are still required as input to our method, and they are used to perform the projection. Figure 2(bottom left) shows an example of a LiDAR-image.
@@
The architecture of CMRNext is based on RAFT [41], a state-of-the-art network for optical flow estimation between two consecutive RGB frames. More details regarding the network architecture are reported in Section III-C. The output of CMRNext is a dense feature map, which consists of four channels that represent, for every pixel in the LiDAR-image, the displacement (, ) of the pixel in the RGB image from the same world point and the uncertainties of this match (, ). A visual representation of this pixel displacement is depicted in Figure 2. The predicted uncertainties can be used for any downstream task, such as active SLAM or map change detection.
@@ III-C Network Architecture
where is the concatenation of the patch from the correlation volume, the output of the context encoder, and the current predicted flow . The operation denotes a convolutional layer with a kernel size of , and are the weights of the convolutional layers, and is the sigmoid operation. Afterward, the hidden state is used to predict the next residual flow and the respective uncertainty using two convolutional layers. The residual flow is then added to the previous flow to obtain the updated flow . This process is repeated times to obtain the final flow. Since the GRU weights and are shared across iterations, the number of update iterations can be increased during inference for improved performance. Finally, the predicted flow at every iteration is upsampled to the original image resolution using learned convex upsampling [50].
@@ IV-G Monocular Localization
Initial Pose | Ground Truth | CMRNext (ours) |
@@
IV-G1 Iterative Refinement
Translation [cm] | Rotation [°] | ||||
Mean | Std | Mean | Std | ||
Stereo | S-PTAM [66] | 871 | 407 | 1.74 | 1.08 |
VINS-Fusion† [67] (w/o LC) | 940 | 4.99 | 2.51 | ||
VINS-Fusion† [67] (w/ LC) | 710 | 2.23 | 3.02 | ||
ORB-SLAM3† [3] (w/o LC) | 328 | 160 | 1.34 | 0.65 | |
ORB-SLAM3† [3] (w/ LC) | 198 | 89 | 1.32 | 0.67 | |
Stereo + Map | Kim et al. [68] | 13 | 11 | 0.32 | 0.39 |
Zuo et al.[69] | 47 | – | 0.87 | – | |
Ding et al. [70] | 45 | – | – | – | |
SemLoc [71] | 34 | – | – | – | |
Monocular + Map | Caselitz et al. [10] | 30 | 11 | 1.65 | 0.91 |
Sun et al. [17] | 11 | 8 | 1.42 | 1.09 | |
Zhang et al. [72] | 58 | – | – | – | |
SemLoc [71] | 149 | – | – | – | |
CMRNet† [24] | 126.02 | 36.67 | |||
I2D-Loc† [65] | 57.76 | 52.41 | |||
I2D-VO† [73] | 24 | 26 | 0.54 | 0.94 | |
I2D-Loc++† [73] | 13 | 17 | 0.49 | 0.55 | |
CMRNet++ [30] | 74 | 35 | 0.42 | 0.25 | |
CMRNext (ours) | 11 | 6 | 0.25 | 0.15 |
LC represents whether the loop closure detection module was enabled (w/) or disabled (w/o). †: Results for these methods are taken from [73].
IV-H Full Trajectory Localization
The experiments reported in Section IV-G assume that an initial pose is provided for every frame of the sequence. However, in real-world scenarios, this initial pose might not always be available, such as in GNSS-denied environments. We therefore perform a trajectory-level evaluation, where we provide an initial pose only for the first frame of the trajectory, and then we initialize the successive frames using the prediction of the previous one. This evaluation protocol includes the unique challenge that the drift accumulated over multiple frames, if not corrected, might yield LiDAR-images that have no overlap with the camera images, making the localization infeasible. We compare the performance of CMRNext with tradition methods for stereo SLAM [66, 67, 3], stereo localization in LiDAR maps [68, 69, 70, 71], monocular localization in LiDAR maps [10, 17, 72, 71], as well as learning-based methods [24, 65, 73, 30]. We compare all methods using the standard Absolute Trajectory Error (ATE) metric typically employed to evaluate odometry and SLAM approaches. The results reported in Table IV show that our CMRNext outperforms all existing methods achieving a mean ATE of and for the translation and rotation components, respectively. Previous methods CMRNet and I2D-Loc suffer from large drift that results in unrecoverable localization failure. I2D-Loc++ explicitly tackles this problem by employing optical flow between consecutive images and multi-frame bundle adjustment optimization, drastically reducing the localization errors. While our method does not employ any tracking or multi-frame processing, it outperforms I2D-Loc++ and notably achieves the lowest standard deviations among all methods due to its robust camera-LiDAR matching capability. In Figure 7, we report the predicted path and the translation errors of CMRNext compared with S-PTAM, ORB-SLAM3, and CMRNet++ on the sequence 00 of the KITTI dataset.
IV-I Ablation Study
In Table V we evaluate the performance of CMRNext using different optical flow networks. In particular, we evaluate established optical flow networks, namely PWCNet [41], MaskFlowNet [74] and RAFT [50], as well as the recent state-of-the-art methods GMFlow [52] and FlowFormer [75]. Unsurprisingly, RAFT outperforms previous methods PWCNet and MaskFlowNet by a large margin. However, we noticed that both transformer-based methods GMFlow and FlowFormer underperformed compared to RAFT, although they achieved better results on optical flow benchmarks. We hypothesize that this is due to the fact that networks based on Vision Transformers (ViT) lack intrinsic inductive bias [76] and thus require a large amount of data to train. The vast majority of methods for optical flow estimation are pretrained on large-scale synthetic datasets, such as FlyingChairs and FlyingThings, and therefore ViT-based methods are able to learn this inductive bias. For our cross-modal task, however, such large-scale datasets are not available, and therefore, the CNN-based method RAFT achieves the best performance among all the evaluated networks. Moreover, the positional encoding used in transformers is sensitive to the shape of the input, and during inference, images with the same resolution as the ones used during training should be used for optimal performance. The authors of FlowFormer [75] suggested a tiling strategy where the input image is split into multiple tiles, each tile is then independently processed by the network, and the resulting optical flow fields are combined together. While this strategy substantially improves the performance of FlowFormer (see Table V), it requires multiple forward passes of the network, drastically increasing the computational cost.
@@ IV-I Ablation Study
In Table VI, we additionally investigate the impact of the positional encoding presented in Equation 10, the occlusion filter discussed in Section III-A, the voxel size used to discretize the map, and which input to use for the context encoder (camera image or LiDAR image). We first evaluate the effect of the positional encoding by comparing the performance of CMRNext using no positional encoding, the positional encoding presented in Equation 10 with six Fourier frequencies, and with 12 frequencies. The first set of rows in Table VI shows that the positional encoding substantially improves the performance of CMRNext, with the best results achieved using 12 frequencies, demonstrating that Fourier features are better suited for deep neural network processing than (possibly normalized) coordinates.
@@
Subsequently, we evaluated three different resolutions for the LiDAR maps: , , and . As shown in the third set of rows in Table VI, the voxel size has a significant impact on the performance of our method. This result is not surprising, as a coarse resolution would lead to sparse LiDAR-images with non-well-defined edges, hindering the pixel-wise matching with the camera images. Although a resolution lower than could further improve the performance, computational and storage requirements would increase significantly. Therefore, we chose as the default resolution for the LiDAR maps, as a good trade-off between performance and computational cost.
Finally, we compared the performance of CMRNext using the camera image and the LiDAR image as the input for the context encoder. The last two rows of Table VI show that the LiDAR context encoder achieves better performance than the camera context encoder.
IV-J LiDAR-Camera Calibration
@@
We compared the performance of CMRNext on the LiDAR-camera extrinsic calibration task with our previous methods CMRNet [24], as well as with the state-of-the-art methods RGGNet [39] and LCCNet [40]. For the latter methods, we used the pretrained models provided by the respective authors444https://github.com/LvXudong-HIT/LCCNet
https://github.com/KleinYuan/RGGNet. Other existing methods [64, 38, 45, 62] do not provide a publicly available implementation and therefore we were unable to evaluate them on the right camera of the KITTI dataset, as discussed above. In Table VII, we report the results of CMRNet, LCCNet, and CMRNext after the iterative refinement (five iterations for CMRNet and LCCNet, and three iterations for CMRNext). The results show that although LCCNet achieves the lowest translation error when evaluated on the same camera used for training, it fails to generalize to a different camera, still predicting the extrinsic for the left camera when evaluated on the right one. We observe a similar behavior for our previous work CMRNet, as both CMRNet and LCCNet are camera-dependent and thus more prone to overfitting to the specific extrinsic. Our proposed CMRNext, on the other hand, outperforms existing methods on the right camera, with a median translation error of and a median rotation error of , while at the same time achieving the lowest rotation error and comparable translation error on the left camera.
@@
where is the initial error applied to the -th sample, and is the left-invariant Riemannian metric on the SE(3) Lie group between the ground truth and the predicted extrinsic calibration. The results reported in Table VIII show that CMRNext outperforms both -RegNet and RGGNet on all testing splits by a large margin. When evaluated on the right camera, CMRNext achieves an MMR of 90.13% on the T1 testing split, compared to 12.37% achieved by RGGNet and 89.71% MRR on the T3 split, compared to -113.67% achieved by RGGNet.
@@
IV-K Extrinsic Calibration — Temporal Aggregation
@@
The multi-frames results reported in Table IX show that the performance of all tendency measures decreases as the number of frames used for aggregation decreases, as hypothesized. Moreover, the mode consistently outperforms the mean measures in all experiments. The median aggregation achieves the best translation error on the left camera but decreases significantly on the right camera, where the mode achieves the best results. When using all frames of the testing sequence (4541 frames), the aggregated mode extrinsic calibration achieves a translation error of and a rotation error of on the left camera, and and on the right camera, setting the new state of the art for targetless LiDAR-camera extrinsic calibration.
IV-L Extrinsic Calibration — Generalization
We evaluate the generalization ability of CMRNext on the extrinsic calibration task on our three in-house platforms detailed in Section IV-A: a self-driving perception car, a quadruped robot, and a UAV. To quantitatively evaluate our proposed method, we first calibrated the extrinsic parameters of each platform using a simple manual calibration procedure. In particular, we recorded a calibration sequence by moving a planar calibration board in front of the platform and then we manually selected the corners of the calibration board in both the camera images and the LiDAR scans. Finally, based on these correspondences, we used PnP with RANSAC, followed by a Levenberg-Marquardt optimization based on the inliers of RANSAC. We compare the translation and rotation errors of CMRNext on the three platforms against the manual calibration. In this experiment, since the number of frames in each dataset is relatively low (around 20 frames), we use 30 different random initial noise poses for each frame, and we report the median frame-wise errors and the error of the pose computed using temporal aggregation across all frames. We compare our approach with the optimization-based method proposed by Koide et al. [79], which leverages temporal aggregation of point clouds and multi-frame optimization. The results reported in Table X show that our method with temporal aggregation achieves rotation errors lower than on all three platforms and translation errors lower than on the car and UAV platforms, outperforming Koide et al. [79] on almost all metrics. The translation error on the quadruped is considerably higher than those on the other platforms. However, this higher error might derive from an inaccurate manual calibration, as we have no way to guarantee the accuracy of our manual calibration procedure.